OpenStudy (anonymous):

Which of the following are solutions to the quadratic equation? Check all that apply. x^2+10x+25=7 Thanks!

OpenStudy (solomonzelman):

\[x^2+10x+25=7\]\[(x+5)^{2}=7\]\[x+5=±\sqrt{7}\]\[x=-5±\sqrt{7}\]

OpenStudy (solomonzelman):

Completing the square, that's what you are working on in class, right ?

OpenStudy (anonymous):

Yes, thank you so much! It's all confusing to me still /: Thank you though!(: